OpenStudy (goformit100):

In 100 m race, A covers the distance in 36 seconds and B in 45 seconds. In this race A beats B by: A. 20 m B. 25 m C. 22.5 m D. 9 m

OpenStudy (goformit100):

@e.cociuba

OpenStudy (goformit100):

@msingh

OpenStudy (goformit100):

@triciasotaso

OpenStudy (goformit100):

@Prudence

OpenStudy (anonymous):

the average speed for both runners is \[v_a =\frac{ 100 }{ 36 }=\frac{ 25 }{ 9 } ,v_b=\frac{ 100 }{ 45 }=\frac{ 20 }{ 9 }\] \[v_a-v_b=\frac{ 5 }{ 9 }\] so\[x=(v_a-v_b)t \implies \frac{ 5 }{ 9 }t\]

OpenStudy (anonymous):

the race finish after t=45 then x=25

OpenStudy (anonymous):

not sure about the logic

OpenStudy (goformit100):

Thanks

OpenStudy (anonymous):

the physical meaning of v_a-v_b can be considereed as the speed needed to cathc up the fastest runner

OpenStudy (anonymous):

yw

OpenStudy (agent0smith):

Find B's velocity, then find how far he's gone after 36 seconds (this is when A crosses the line, ie he's run 100m). Subtract that distance from 100m (eg if B has only run 75m, he's 25m behind when A finishes)

OpenStudy (kropot72):

You need to know what distance B had travelled at the moment when A finished the 100 m. This distance cannot be found from the information given. For example B could possibly have been only 10 m behind when A crossed the finishing line and then slowed down drastically over the last 10 m.

OpenStudy (agent0smith):

Good point @kropot72 but I think it's reasonable to assume an average velocity.

OpenStudy (agent0smith):

Although, given that it took FORTY FIVE SECONDS to finish a 100m race... you might be right.

OpenStudy (kropot72):

@agent0smith You are right, that was part of my thinking on this one :)